### v1

@@ IV Solving the BDKG field equations
IV.0.1 Solving for the scalar fields
We shall employ the following identities
@@ IV.0.1 Solving for the scalar fields
$=\tilde{g}^{\mu\nu}\tilde{\nabla}_{\nu}(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$ (79)
$=\tilde{g}^{\mu\nu}\left[\partial_{\nu}(e^{-\theta}\partial_{\mu}\Psi)-\tilde{\Gamma}_{\mu\nu}^{\lambda}e^{-\theta}\partial_{\lambda}\Psi\right]$ (80)
$=\tilde{g}^{\mu\nu}\partial_{\nu}(e^{-\theta}\partial_{\mu}\Psi)-\left(\tilde{g}^{\mu\nu}\tilde{\Gamma}_{\mu\nu}^{\lambda}\right)e^{-\theta}\partial_{\lambda}\Psi$ (81)
@@
in which the identity (77) has been employed. As $\theta$ and $\Psi$ are functions of $u$, the second term imposes $\lambda=u$, leaving
$\partial_{\mu}(\sqrt{-\tilde{g}}\,\tilde{g}^{u\mu})=\partial_{u}(\sqrt{-\tilde{g}}\,\tilde{g}^{uu})=0$ (83)
The left-hand-side of Eq. (73) becomes
~$\tilde{\nabla}^{\mu}(e^{-\theta}\tilde{\nabla}_{\mu}\Psi)$ $=\tilde{g}^{uu}\partial_{u}(e^{-\theta}\partial_{u}\Psi)$ (84)
@@
which then yields
@@
with $\Psi_{*}$ being an integration constant, and $\phi$ the BD scalar field in Jordan frame. With the aid of (86), Eq. (72) yields
$(2\omega+3)\,e^{-\theta}\tilde{\square}\,\theta$ $=-\tilde{g}^{uu}\left(e^{-\theta}\Psi^{\prime}\right)^{2}$ (87)
$\Longrightarrow\ (2\omega+3)\,e^{-\theta}e^{2\gamma-2\alpha}\theta^{\prime\prime}$ $=-e^{2\gamma-2\alpha}\Psi_{*}^{2}$ (88)
or
@@
The case of $\Psi_{*}=0$ would correspond to the known vacuum BD solution. However, we find it more convenient to keep $\Psi_{*}\neq 0$ and manage to recover the BD case in the different manner, to be clarified momentarily. Equation (90) has a solution
@@
IV.0.2 Recasting the scalar fields using the limit of $\Psi_{*}\rightarrow 0$
@@
or
$e^{-2\sqrt{\kappa}u_{0}}=\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}$ (96)
bringing Eq. (94) to
@@
$\frac{-2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\sinh^{2}\sqrt{\kappa}(u+u_{0})}$ $=\frac{-8(2\omega+3)\,\kappa}{\Psi_{*}^{2}e^{2\sqrt{\kappa}u_{0}}}\frac{e^{-2\sqrt{\kappa}u}}{\left(1-e^{-2\sqrt{\kappa}(u+u_{0})}\right)^{2}}$ (98)
$=\frac{\phi_{0}\,e^{-2\sqrt{\kappa}u}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)^{2}}$ (99)
@@
IV.0.3 Solving for the metric components
@@
$\alpha^{\prime\prime}$ $=2e^{\alpha}$ (108)
$\Longrightarrow\ \ \alpha(u)$ $=\frac{\chi}{\sinh^{2}\sqrt{\chi}\,u}\,.$ (109)
@@
$\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$
$=\left(\omega+3/2\right)\theta^{\prime 2}+e^{-\theta}\Psi^{\prime 2}$ (113)
$=\left(\omega+3/2\right)\left(-\frac{2\Psi_{*}^{2}}{2\omega+3}e^{\theta}+4\kappa\right)+e^{-\theta}(\Psi_{*}e^{\theta})^{2}$ (114)
$=2(2\omega+3)\,\kappa$ (115)
@@
IV.0.4 Removal of $u_{1}$
The parameter $u_{1}$ in (117) can be removed via rescaling of variables and parameters, given by
@@ IV.0.4 Removal of $u_{1}$
Upon the rescaling, the parameter $u_{1}$ in Eq. (117) is removed, whereas Eqs. (101), (103), and the “constraint” (116) are unaffected.
V Exact solution of BDKG action in the Jordan frame
@@
The two scalar fields expressed above are obviously related per Eq. (86):
$\frac{d\Psi(u)}{du}=\Psi_{*}\,\phi(u)\,.$ (124)
The integration constants entering this solution are $\{a,\,\chi,\,\phi_{0},\,\Psi_{*}\}$.
@@
The final solution (121)–(123) can also be expressed in a more familiar coordinate. Define $r:=1/u$. The limit $u\rightarrow 0$ would correspond to the spatial infinity $r\rightarrow\infty$. The metric (121) becomes
@@
VI Behaviors at large $\omega$
@@
First of all, the static spherisymmetric metric for BDKG theory is conformally equivalent to the FJNW spacetime
@@
For a given pair of $\{\chi,\,a\}$, the metric of BDKG gravity and the FJNW metric would share a common causal structure. For a given casual structure, viz. keeping $\chi$ and $a$ fixed, the dependence on $\omega$ is encapsulated in the conformal factor, viz. $1/\phi(r)$.
VI.0.1 The $\omega\rightarrow\infty$ behavior of the scalar fields
@@
$-\frac{\phi_{0}\sqrt{2(\chi-a^{2})}\left(1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{3}}\frac{1}{\sqrt{\omega\,}r}+\mathcal{O}\left(\frac{1}{\omega}\right)$ (129)
That is to say, apart from the constant term $\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$, the subleading contribution to the $\phi$ field is in the order of $1/\sqrt{\omega}$.
Similarly, the Taylor expansion of Eq. (127) yields
@@
VI.0.2 The $\omega\rightarrow\infty$ “remnant” metric
With $\chi$ and $a$ held fixed, $\kappa$ behaves as $1/\sqrt{|\omega|}$. Upon sending $\omega$ to infinity, the BD scalar field thus becomes a constant field. However, the limiting metric is not necessarily a Schwarzschild metric. In the harmonic radial coordinate
@@ VI.0.2 The $\omega\rightarrow\infty$ “remnant” metric
whereas $\phi(u)\equiv\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$ and $\Psi(u)\equiv\text{const}$. Remark: a constant $\Psi$ has no physical impact at all (while a constant $\phi$ is physically present in the field equations). This is the metric part of the FJNW solution of GR coupled with a scalar field in the form
$S_{\text{BD}}=\int d^{4}x\sqrt{-g}\,\phi_{0}\biggl{(}\mathcal{R}-g^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi\biggr{)}$ (132)
a point that we will justify in the next section.
VII Formal proof of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly
In [25, 24] one of the present authors introduced a set of ideas aimed at establishing the ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior for matter with non-vanishing EMT trace. This section builds upon those ideas, working towards a comprehensive proof. Departing from the conventional expansion in terms of $1/\omega$, we initiate our proof with the following expansions:
@@ VII Formal proof of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly
VII.1 Proof for BDKG theory
Let us insert the expansions (133)–(135) into the original action (21). To the leading order, we obtain
$\mathcal{S}_{\text{BDKG}}=\ \stackrel{{\scriptstyle 0}}{{\mathcal{S}}}+\mathcal{O}\left(\frac{1}{\sqrt{\omega}}\right)$ (136)
in which the leading term reads
[…]
The action $\stackrel{{\scriptstyle 0}}{{\mathcal{S}}}$ is effectively a “remnant” of the BDKG action upon sending $\omega$ to infinity. The tensor $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ represents a “background” metric. It effectively describes GR in presence of $\stackrel{{\scriptstyle 0}}{{\Psi}}$, but augmented by the scalar $\stackrel{{\scriptstyle 1}}{{\varphi}}$, which is remnant from the vanishing part of the BD scalar. The term $\bar{\phi}^{-1}$ assumes the role of a Newtonian gravitational constant.
Upon variation, the “remnant” action (137) produces the field equations
$\stackrel{{\scriptstyle 0}}{{\mathcal{R}}}_{\mu\nu}$ $=\partial_{\mu}\stackrel{{\scriptstyle 1}}{{\varphi}}\partial_{\nu}\stackrel{{\scriptstyle 1}}{{\varphi}}+\bar{\phi}^{-1}\partial_{\mu}\stackrel{{\scriptstyle 0}}{{\Psi}}\partial_{\nu}\stackrel{{\scriptstyle 0}}{{\Psi}}$ (138)
$\stackrel{{\scriptstyle 0}}{{\square}}\ \stackrel{{\scriptstyle 0}}{{\Psi}}$ $=0$ (140)
Here, the superscript 00 indicates terms and operators evaluated with respect to the background metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$. If a solution to these equations exist for a given $\stackrel{{\scriptstyle 0}}{{\Psi}}$, then this would establish two things:
  1. 1.
    The two non-vanishing terms $\partial_{\mu}\stackrel{{\scriptstyle 0}}{{\Psi}}\partial_{\nu}\stackrel{{\scriptstyle 0}}{{\Psi}}$ and $\partial_{\mu}\stackrel{{\scriptstyle 1}}{{\varphi}}\partial_{\nu}\stackrel{{\scriptstyle 1}}{{\varphi}}$ have independent contributions. Whereas the contribution from $\stackrel{{\scriptstyle 0}}{{\Psi}}$ is expected as a result of the matter source $\Psi$, the contribution from $\stackrel{{\scriptstyle 1}}{{\varphi}}$ is emergent as a result of the BD field dynamics. The latter contribution acts as an additional source to the “background” metric $\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}$ which then must deviate from a GR solution in the presence of only the KG matter source $\Psi$; and
  2. 2.
    The $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature in the BD scalar field is self-transparent via $\stackrel{{\scriptstyle 1}}{{\varphi}}$, per Eq. (133).
Specifying to static spherisymmetric setup, it is straightforward to show that the field equations above can be interpreted as an extension of FJNW to two independent scalar fields to have  [^8] We are using the coordinate $r$ which is a reciprocal of the harmonic radial coordinate $u$, viz. $r:=1/u$.
$\stackrel{{\scriptstyle 1}}{{\varphi}}$ $\ =\frac{\sigma}{r}$ (141)
$\stackrel{{\scriptstyle 0}}{{\Psi}}$ $\ =\frac{\Sigma}{r}$ (142)
in which $\sigma$ and $\Sigma$ are “scalar charges”, a term coined by Bronnikov [26], and the “background” metric is given by
$\stackrel{{\scriptstyle 0}}{{g}}_{\mu\nu}dx^{\mu}dx^{\nu}$ $=-e^{-\frac{2a}{r}}dt^{2}$
$+e^{\frac{2a}{r}}\biggl{(}\frac{(\sqrt{\chi}/r)^{4}}{\sinh^{4}\sqrt{\chi}/r}dr^{2}+\frac{(\sqrt{\chi}/r)^{2}}{\sinh^{2}\sqrt{\chi}/r}\,r^{2}d\Omega^{2}\biggr{)}$ (143)
in which $\chi$, $a$, $\sigma$ and $\Sigma$ are related by
$2\chi=2a^{2}+\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$ (144)
Due to the presence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$, the total scalar charge squared gets modified, from $\bar{\phi}^{-1}\Sigma^{2}$ to $\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$, affecting the FJNW metric. The BD scalar field $\phi$ thus acquires a behavior
$\phi=\bar{\phi}\,\Bigl{(}1+\frac{\sigma}{\sqrt{\omega}\,r}+\dots\Bigr{)}$ (145)
This completes our proof. When $\sigma=0$ we recover the limiting GR with the original massless Klein-Gordon field as its only source.
VII.2 Verification
The result of limiting $\omega\rightarrow\infty$ has been obtained without resorting to the exact solution. It is desirable to confirm the result using the exact solution. From (129) we extract
$\bar{\phi}\equiv\frac{\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}$ (146)
and
$\stackrel{{\scriptstyle 1}}{{\varphi}}\ =-\frac{1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}\frac{\sqrt{2(\chi-a^{2})}}{r}$ (147)
Similarly, from (130)
$\stackrel{{\scriptstyle 0}}{{\Psi}}\ =\frac{\Psi_{*}\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}\frac{1}{r}$ (148)
the function form of which is consistent with (141) and (142). The scalar charges are
$\sigma$ $=-\frac{1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}{1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}}\sqrt{2(\chi-a^{2})}$ (149)
$\Sigma$ $=\frac{\Psi_{*}\phi_{0}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}}$ (150)
Then, from (146), (149) and (150), it is straightforward to verify that
$\sigma^{2}+\bar{\phi}^{-1}\Sigma^{2}$ $=2(\chi-a^{2})$ (151)
in perfect agreement with (144). Hence, the exact solution is compatible with the proof produced in the preceding subsection.
VII.3 Generalizing to a generic matter source
The proof presented in Section VII.1 can be extended beyond the massless KG field, however. Let us consider a BD action coupled with a matter field $\Psi$ in the form
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +L^{(m)}\bigl{(}\{g_{\alpha\beta},\phi\};\Psi\bigr{)}\Bigr{]}$ (152)
The matter sector can be of generic nature, scalar, spinor, massive, etc. It may also couple with the BD scalar field and/or couple with the metric tensor in a non-minimal fashion.
@@ VII.3 Generalizing to a generic matter source
For the static spherisymmetric setup, this statement can be proved. The harmonic equation for $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in the harmonic radial coordinate yields a solution
@@
with $\sigma$ being a scalar charge  [^10] It should be commented that in the harmonic radial coordinate, the spatial infinity corresponds to $u=0$. The divergence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in Eq. (158) when $u$ goes to infinity means that the solution is diverging at a finite $r$. Thence the possibility of the non–GR behavior, in this static case, is related to the presence of a singularity at a finite $r$.. The non-constant $\stackrel{{\scriptstyle 1}}{{\varphi}}$ thus adds a constant contribution to the $uu-$equation, per
@@
VIII Conclusion and outlook

### v3

@@ III.3 Solving for the scalar fields
The following identities will be useful
@@
in which the identity (45) has been employed. Equation (41) then yields
@@
with $\Psi_{*}$ being an integration constant. With the aid of (49), Eq. (40) yields
@@
which has a solution
@@
Recasting the scalar fields using the limit of $\Psi_{*}\rightarrow 0$:
@@
and bring Eq. (55) to
@@
$\frac{-2(2\omega+3)\,\kappa}{\Psi_{*}^{2}\,\sinh^{2}\sqrt{\kappa}(u+u_{0})}$ $=\frac{\phi_{0}\,e^{-2\sqrt{\kappa}u}}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(2\omega+3)\kappa}e^{-2\sqrt{\kappa}u}\right)^{2}}$ (58)
@@
III.4 Solving for the metric components
@@
$\alpha^{\prime\prime}=2e^{\alpha}\ \ \Longrightarrow\ \ \alpha(u)$ $=\frac{\chi}{\sinh^{2}\sqrt{\chi}\,u}\,.$ (67)
@@
$\left(\omega+3/2\right)\partial_{\mu}\theta\partial_{\nu}\theta+e^{-\theta}\partial_{\mu}\Psi\partial_{\nu}\Psi$ $=2(2\omega+3)\,\kappa$ (71)
@@
Removal of $u_{1}$:
The parameter $u_{1}$ in (73) can be removed via rescaling of variables and parameters, per
@@
Let us remark that Eqs. (60), (62), and the “constraint” (72) are unaffected.
III.5 Exact solution of BDKG action in Jordan frame
@@
The two scalar fields expressed above are obviously related, in accordance with Eq. (49). The integration constants entering this solution are $\{a,\,\chi,\,\phi_{0},\,\Psi_{*}\}$.
@@
The final solution (77)–(79) can also be expressed in a more familiar coordinate. Define
$r:=1/u\,.$ (80)
The limit $u\rightarrow 0$ would correspond to the spatial infinity $r\rightarrow\infty$. The metric (77) becomes
@@
III.6 The $\mathcal{O}\left(1/\sqrt{\omega}\right)$ anomaly in BDKG
@@
First of all, the static spherically symmetric metric for BDKG theory is conformally equivalent to the FJNW spacetime
@@
For a given pair of $\{\chi,\,a\}$, the metric of BDKG gravity and the FJNW metric would share a common causal structure. For a given causal structure, viz. keeping $\chi$ and $a$ fixed, the dependence on $\omega$ is encapsulated in the conformal factor, viz. $1/\phi(r)$.
III.6.1 The $\omega\rightarrow\infty$ behavior of the scalar fields
@@
$-\frac{\phi_{0}\sqrt{2(\chi-a^{2})}\left(1-\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)}{\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{3}}\frac{1}{\sqrt{\omega\,}r}+\mathcal{O}\left(\frac{1}{\omega}\right).$ (85)
That is to say, apart from the constant term $\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$, the sub-leading contribution to the $\phi$ field is in the order of $1/\sqrt{\omega}$. Similarly, the Taylor expansion of Eq. (83) yields
@@
III.6.2 The $\omega\rightarrow\infty$ “remnant” metric
With $\chi$ and $a$ held fixed, $\kappa$ behaves as $1/\sqrt{|\omega|}$. Upon sending $\omega$ to infinity, the BD scalar field thus becomes a constant field. However, the limiting metric is not a Schwarzschild metric, but an FJNW one. In the harmonic radial coordinate
@@ III.6.2 The $\omega\rightarrow\infty$ “remnant” metric
whereas $\phi(u)\equiv\bar{\phi}:=\phi_{0}/\left(1+\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}\right)^{2}$ and $\Psi(u)\equiv\text{const}$. Remark: a constant $\Psi$ has no physical impact at all (while a constant $\phi$ is physically present in the field equations). Expression (87) is the metric part of the FJNW solution of GR sourced by a scalar field
$S_{\text{BD}}=\int d^{4}x\sqrt{-g}\,\phi_{0}\biggl{(}\mathcal{R}-g^{\mu\nu}\nabla_{\mu}\varphi\nabla_{\nu}\varphi\biggr{)}\,.$ (88)
IV Formal proof of the ${\cal O}\left(1/\sqrt{\omega}\right)$ anomaly for Brans–Dicke action sourced by any matter field
In this section, we consider a BD action coupled with a matter field $\Psi$ in the form
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +L^{(m)}\bigl{(}\{g_{\alpha\beta},\phi\};\Psi\bigr{)}\Bigr{]}$ (89)
The matter sector can be of generic nature; the $\Psi$ field can be scalar, spinor, massive, and so forth. It may also couple with the BD scalar field and/or couple with the metric tensor in a non-minimal fashion. The EMT of the matter sector can have a non-vanishing trace.
In [25, 24] one of the present authors (BC) introduced a set of ideas aimed at establishing the ${\cal O}\left(1/\sqrt{\omega}\right)$ behavior for matter with non-vanishing EMT trace, such as in the action (89). This section builds upon those ideas, working towards a comprehensive proof. Departing from the conventional expansion in terms of $1/\omega$, we initiate our proof, as made in [25] for the scalar and the metric, with the following expansions:
@@
For the static spherically symmetric setup, this statement can be proved. The harmonic equation for $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in the harmonic radial coordinate yields a solution
@@
with $\sigma$ being a scalar charge. It should be commented that in the harmonic radial coordinate, as per Eq. (80), the spatial infinity corresponds to $u=0$. The divergence of $\stackrel{{\scriptstyle 1}}{{\varphi}}$ in Eq. (98) when $u$ goes to infinity means that the solution is diverging at $r=0$. Consequently, the possibility of the non–GR behavior, in this static case, is related to the presence of a singularity at $r=0$.
The non-constant $\stackrel{{\scriptstyle 1}}{{\varphi}}$ thus adds a constant contribution to the $uu-$equation, per
@@
V The Robertson parameters
In this section, we derive the Robertson (or Eddington-Robertson-Schiff) parameters, $\beta$ and $\gamma$, which are physical, i.e. measurable, quantities, from our exact BDKG solution.
For GR, it is known that $\beta_{\,\text{GR}}=\gamma_{\,\text{GR}}=1$. The Robertson expansion in isotropic coordinates is, using the relativistic units [8]:
$ds^{2}$ $=-\left(1-2\,\frac{M}{\rho}+2\beta\,\frac{M^{2}}{\rho^{2}}+\dots\right)dt^{2}$
$\ \ \ \,+\left(1+2\gamma\,\frac{M}{\rho}+\dots\right)\left(d\rho^{2}+\rho^{2}d\Omega^{2}\right)$ (102)
in which $\beta$ and $\gamma$ are the Robertson parameters. The metric of Eq. (127) in Appendix B yields
$\phi_{0}\,ds^{2}$ $=-\biggl{(}1-\frac{4K\left[(a+\sigma)L+(a-\sigma)\right]}{\sqrt{\chi}\left(1+L\right)\rho}$
$+\frac{8K^{2}\left[(a+\sigma)^{2}L^{2}+2a^{2}L+(a-\sigma)^{2}\right]}{\chi\,(1+L)^{2}\rho^{2}}+\dots\biggr{)}$
×$\times\left(1+L\right)^{2}dt^{2}$
$+\biggl{(}1+\frac{4K\left[(a-\sigma)L+(a+\sigma)\right]}{\sqrt{\chi}\left(1+L\right)\rho}+\dots\biggr{)}$
×$\times\left(1+L\right)^{2}\left(d\rho^{2}+\rho^{2}d\Omega^{2}\right)$ (103)
where $L:=\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}$. Upon absorbing $(1+L)$ into $dt$, $d\rho$ and $\rho$, and $\phi_{0}$ into $ds^{2}$ in Eq. (103) and comparing it with Eq. (102), we have
$M$ $=\frac{2K}{\sqrt{\chi}}\left[(a+\sigma)L+(a-\sigma)\right]$ (104)
$\beta$ $=1+\frac{2\sigma^{2}L}{\left[(a+\sigma)L+(a-\sigma)\right]^{2}}$ (105)
$\gamma$ $=1-\frac{2\sigma\left(L-1\right)}{(a+\sigma)L+(a-\sigma)}$ (106)
Employing the relation (128), the Robertson parameters become
$\beta$ $=1+\frac{\Psi_{*}^{2}\phi_{0}/(4a^{2})}{\Bigl{[}\Bigl{(}\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}+1\Bigr{)}\sqrt{2\omega+3}+\Bigl{(}\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}-1\Bigr{)}\frac{\sqrt{\chi-a^{2}}}{a}\Bigr{]}^{2}}$ (107)
$\gamma$ $=1-\frac{2\left(\frac{\Psi_{*}^{2}\phi_{0}}{8\left(\chi-a^{2}\right)}-1\right)\frac{\sqrt{\chi-a^{2}}}{a}}{\Bigl{(}\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}+1\Bigr{)}\sqrt{2\omega+3}+\Bigl{(}\frac{\Psi_{*}^{2}\phi_{0}}{8(\chi-a^{2})}-1\Bigr{)}\frac{\sqrt{\chi-a^{2}}}{a}}$ (108)
These formulae evidently recover $\beta_{\,\text{GR}}=\gamma_{\,\text{GR}}=1$ in the infinite $\omega$ limit, and exhibit behavior on order of $\mathcal{O}\left(1/\sqrt{\omega}\right)$ for a fixed set of $a$ and $\chi$. As physical quantities, $\beta$ measures the degree of nonlinearity in superposition of gravity, while $\gamma$ specifies the amount of spatial curvature produced by the mass source [39, 40]. Therefore, the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ signature in $\beta$ and $\gamma$ carry meaningful physical implications in BD configurations involving trace-carrying KG matter fields.
VI Physical implications
Besides the $\mathcal{O}\left(1/\sqrt{\omega}\right)$ imprints in the Roberson parameters $\beta$ and $\gamma$ as presented in the preceding section, let us further stress that the issue tackled in this paper is not just a question of formal mathematics, but that genuine potentially new physical features are at stake, in both the EMT trace $T=0$ and $T\neq 0$ cases. Indeed, it results from an $\mathcal{O}\left(1/\sqrt{\omega}\right)$ behavior of the BD scalar field that the large $\omega$ Brans-Dicke vacuum spherical solution (i.e. Brans class I if $\omega>-3/2$) does not converge to the GR vacuum spherical solution (i.e. Schwarzschild), but to the GR spherical solution filled by a massless scalar (i.e. the FJNW solution). Especially in the strongly scalarized case, the FJNW solution exhibits drastically specific features with respect to Schwarzschild. Indeed, for a given gravitational mass of the field, circular orbits exist for any areal radius (while the areal radius is lower bounded in Schwarzschild), with a diverging orbital frequency measured by a remote observer (while this frequency is upper bounded in Schwarzschild), despite the Einstein gravitational red-shifting [41]. Since the associated gravitational wave (GW) frequency, in the Extreme Mass Ration Inspiral (EMRI) scheme, is twice the orbital one, the potential impact of the effect in the new GW astronomy is obvious. Still in the strongly scalarized case, another qualitative specific features with respect to Schwarzschild concerns the light bending, which experiments a maxima when the impact parameter decreases. This results in the presence of a caustic surface, which yields an infinite amplification of the apparent luminosity of a pointlike star crossing this caustics [42]. Thence, the case of an orbiting star on a properly inclined orbit results in a sequence of luminosity peaks, an effect which does not exist in the Schwarzschild spacetime. In the $T\neq 0$ case, the age of the Universe is not constrained to a fixed value by the present Hubble parameter $H_{0}$ in BD dust filled flat Robertson-Walker cosmology, even in the $\omega\rightarrow\infty$ limit, while it is fixed to $2/(3H_{0})$ in GR dust filled flat Robertson-Walker cosmology [24]. We plan to visit this issue in the future.
VII Conclusion and outlook